OpenStudy (anonymous):

which is a solution to the equation y = 5x-6?

OpenStudy (texaschic101):

do you have answer choices ?

OpenStudy (anonymous):

A) (3,21) B) (4, 2) C) (5, 35) D) (6, 24) which one

OpenStudy (texaschic101):

that would be D...(6,24) y = 5x - 6 24 = 5(6) - 6 24 = 30 - 6 24 = 24 (correct)

OpenStudy (texaschic101):

I just subbed in my points (x,y) into the equation to see which one made the equation true